OpenStudy (anonymous):

what is the sq rt of 30/x^4 ? Is it numerator sq rt 30x^4 over denominator x^4 ?

OpenStudy (anonymous):

depends on the notation: is it \[\sqrt{\frac{ 30 }{ x^{4} }}\] or \[\frac{ \sqrt{30} }{ x^{4} }\]

OpenStudy (anonymous):

\[\sqrt{\frac{ 30 }{ x^4 }}\]

OpenStudy (anonymous):

I think the answer is \[\sqrt{\frac{ 30x^4}{ x^4 }}\]

OpenStudy (anonymous):

sorry i meant i think the answer is \[\frac{ \sqrt{30x^4} }{ x^4}\]

OpenStudy (anonymous):

take the square root of \[x^{4}\] of out the square root; \[x ^{2}\frac{ \sqrt{30} }{ x^{4} } \] then subtract exponential to end up with \[\frac{ \sqrt{30} }{ x^2 }\]

OpenStudy (anonymous):

Thank you very very much.